Grade 128 grade maths

Using identities, evaluate.
102^2 ???????????????????

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

102^2= (100 + 2)^2
= 100^2+ 400 + 2^2
= 10000 + 400 + 4
= 10404

Thanks